OpenStudy (anonymous):

Whats the patter: X Y 4 11 0 -1 3 8 10 ? ? 5 *Hint* Multiply by and number then subtract

OpenStudy (anonymous):

lets say Y=K*X 11=K*4 K=11/4 now when X= 10 K=Y/X or Y=K*X=11/4 * 10 = 55/2 and when Y=5 X=Y/K= 5/11/2= (5*2)/11=10/11

OpenStudy (anonymous):

Thats a little confusing

OpenStudy (anonymous):

a bit of manipulation

OpenStudy (anonymous):

I'd say

OpenStudy (anonymous):

k